# DocKylin: A Large Multimodal Model for Visual Document Understanding with Efficient Visual Slimming

### v1

@@ 3.3 Dynamic Token Slimming
The pseudo-code for the specific implementation can be found in Algorithm 3 in the appendix.
| Stage | Datasets |
| Pre-training | CASIA-HWDB [33], CCpdf [66], ChartQA [40], DeepForm [57], DocVQA [42], DVQA [20], FigureQA [21], InfoVQA [41], KLC [54], OCRVQA [44], PlotQA [43], PubTabNet [79], RVL-CDIP [15], SCUT-HCCDoc [77], SynthDog [22], TAT-QA [80], VisualMRC [59], WildReceipt [56] |
| Instruction tuning | ChartQA [40], DeepForm [57], DocILE [52], DocVQA [42], DVQA [20], InfoVQA [41], InsuranceQA [55], KLC [54], OCR-TROSD[64], TAT-QA [80], VisualMRC [59], WildReceipt [56], WTQ [46] |
4 Experiments
4.1 Implementation Details
Our model underwent two training stages: pre-training and instruction tuning. Training data used in each phase are listed in Table 2. During the pre-training stage, we trained only the visual encoder and the MLP, keeping the LLM parameters fixed. In this phase, the model learns to perform full-text recognition on document images and convert charts and tables into markdown formats [16]. The goal is to enable the visual encoder to comprehend document content and align the visual features with the language space through the MLP.
During the pre-training phase, the model was trained for 450,000 iterations with a batch size of 8. The learning rate decayed from 1e-4 to 1e-5 using cosine annealing, which took approximately 40 A800 GPU-days. In the instruction tuning phase, the model underwent 300,000 iterations with the same batch size and learning rate as in the pre-training phase. This phase further consumed approximately 24 A800 GPU-days. Our Adaptive Pixel Slimming method is employed in both phases, while the Dynamic Token Slimming (DTS) is used only during the instruction tuning phase. The maximum allowed number of pixels, MAX_SIZE, is set to 17281728. We initialize our visual encoder with Donut-Swin (0.07B) [22], pre-trained on the reading texts task, and use Qwen-7B-Chat [2] as our LLM.
4.2 Comparisons on VDU Benchmarks
We first compare our model with several existing general-purpose multimodal large language models (MLLMs). As these models sometimes tend to provide detailed responses, following [78], we adopt accuracy as our metric: a response is considered correct if it contains the complete answer. The results, as shown in Table 3, demonstrate that DocKylin achieves a significant advantage over the competing models. This highlights that VDU tasks remain challenging for current general-purpose MLLMs. Note that descriptions of all these benchmarks are included in the Appendix.
We also compare our model with several current text-centric MLLMs. The results, presented in Table 4, use the same metrics as those employed in the original benchmarks. Detailed descriptions of the specific metrics used for each benchmark can be found in the Appendix. Our method demonstrates advantages over these text-centric MLLMs, achieving leading performance across multiple benchmarks. Notably, our visual encoder is more lightweight than those used by the compared methods.
| Methods | Visual Encoder (Params.) | DocVQAval | InfoVQAval | ChartQA | FUNSD | SROIE | POIE |
| mPLUG-Owl [75] | CLIP-L (0.3B) | 7.4 | 20.0 | 7.9 | 0.2 | 0.1 | 0.3 |
| InstructBLIP [11] | CLIP-G (2B) | 4.5 | 16.4 | 5.3 | 0.2 | 0.6 | 1.0 |
| LLaVA1.5 [35] | CLIP-L (0.3B) | 8.5 | 14.7 | 9.3 | 0.2 | 1.7 | 2.5 |
| Qwen-VL [3] | CLIP-G (2B) | 48.1 | 23.9 | 53.4 | 23.9 | 34.5 | 20.6 |
| Monkey [29] | CLIP-G (2B) | 50.1 | 25.8 | 54.0 | 24.1 | 41.9 | 19.9 |
| InternVL [9] | IViT (6B) | 28.7 | 23.6 | 45.6 | 6.5 | 26.4 | 25.9 |
| InterLM-XComposer2 [12] | CLIP-L (0.3B) | 39.7 | 28.6 | 51.6 | 15.3 | 34.2 | 49.3 |
| DocKylin (ours) | Swin (0.07B) | 65.1 | 34.8 | 58.7 | 25.5 | 49.5 | 36.1 |
| Methods | Visual Encoder (Params.) | DocVQAtest | InfoVQAtest | ChartQA | VisualMRC | DeepForm | WTQ |
| Donut [22] | Swin (0.07B) | 67.5 | 11.6 | 41.8 | 93.9 | 61.6 | 18.8 |
| Pix2Struct[26] | - | 76.6 | 40.0 | 58.6 | - | - | - |
| mPLUG-Doc [75] | CLIP-L (0.3B) | 62.2 | 38.2 | 57.4 | 188.8 | 42.6 | 26.9 |
| UReader [74] | CLIP-L (0.3B) | 65.4 | 42.2 | 65.7 | 221.7 | 49.5 | 29.4 |
| DocPeida [13] | Swin (0.1B) | 47.1 | 15.2 | 46.9 | - | - | - |
| HRVDA [72] | Swin | 72.1 | 43.5 | 67.6 | 211.5 | 63.2 | 31.2 |
| Vary [70] | ViTDet-B & CLIP-L (0.4B) | 76.3 | - | 66.1 | - | - | - |
| TextMonkey [37] | CLIP-G (2B) | 71.5 | - | 65.5 | - | 61.6 | 30.6 |
| TextHawk [76] | SigLIP-SO (0.4B) | 76.4 | 50.6 | 66.6 | - | - | 34.7 |
| LayoutLLM-LLaMA2 [17] | LayoutLMv3-L (0.3B) | 74.3 | - | - | - | - | - |
| DocKylin (ours) | Swin (0.07B) | 77.3 | 46.6 | 66.8 | 319.9 | 64.2 | 32.4 |
@@ 4.2 Comparisons on VDU Benchmarks
| FastV (K=2, R=50%) [8] | ✗ | 306 | 8.4 | 306 | 14.4 | 306 | 2.0 | |
| PruMerge [49] | ✓ | 123 | 5.3 | 123 | 13.8 | 121 | 0.4 | |
| PruMerge+ [49] | ✓ | 232 | 5.2 | 233 | 14.0 | 229 | 0.3 | |
@@
To further demonstrate the superiority of our method, we compared it with other parameter-free visual token compression methods, using LLaVA1.5 [35] as our baseline model (without retraining). The results in Table 5 indicate that our APS and DTS achieve performance improvements with shorter sequence lengths compared to the baseline model, significantly outperforming existing SOTA compression methods. It is worth noting that these results are achieved without retraining. Typically, after retraining, models can better adapt to the new mode of compressed token sequences, leading to further performance improvements [49, 8].
4.3 Ablation Studies and Potential Broader Applications
@@
We validate the effectiveness of APS and DTS on DocKylin and give the results in Table 6. Note that, settings without DTS in Table 6 refer to retraining a new model from scratch without employing DTS in the instruction tuning phase, rather than simply removing DTS from the already trained DocKylin model. The results demonstrate that both APS and DTS effectively reduce token sequence length and enhance computational efficiency, while also improving performance. This aligns with findings from previous studies [37], which suggest that excessively long sequences due to high resolution may degrade performance. This degradation likely occurs because the visual sequences contain too many redundant tokens, placing greater demands on the model’s ability to handle long contexts and making the learning more challenging. Notably, the APS and DTS methods can also be used online during the training phase to accelerate training. Specifically, applying the DTS method during the instruction tuning phase can reduce training time by approximately 15%.
| APS | DTS | Avg. Token Len | Training Time | DocVQAval | |
| Acc. | Latency | ||||
| ✗ | ✗ | 2770 | 1 | 63.1 | 1 |
| ✓ | ✗ | 2198 | 0.97 | 64.5 | 0.81 |
| ✓ | ✓ | 1250 | 0.86 | 65.1 | 0.61 |
| Methods | Resolution | TextVQA | DocVQAval | InfoVQAval | SROIE | FUNSD |
| LLaVA1.5 [35] | 336336 | 38.7 | 8.5 | 14.7 | 1.7 | 0.2 |
| LLaVA1.5 [35] + APS | 336336 | 40.2 (+4.1%) | 10.7 (+27.4%) | 14.7 (+0%) | 3.7 (+118%) | 0.92 (+360%) |
| Qwen-VL [3] | 448448 | 62.1 | 48.1 | 23.9 | 34.5 | 20.6 |
| Qwen-VL [3] + APS | 448448 | 62.4 (+0.4%) | 51.2 (+6.4%) | 24.7 (+4.1%) | 40.0 (+15.9%) | 24.3 (+17.9%) |
| Monkey [29] | 896896 | 64.3 | 50.1 | 25.8 | 41.9 | 24.1 |
| Monkey [29] + APS | 896896 | 65.2 (+1.4%) | 56.3 (+12.4%) | 27.5 (+6.6%) | 47.0 (+12.2%) | 27.3 (+13.3%) |
@@
Since APS is a parameter-free method that requires no additional training or modification of model architecture, it is easily applicable to other MLLMs to enhance their performance on high-resolution document images. To demonstrate its effectiveness, we apply APS to various existing MLLM models (employing it before the image processing module of these models). As shown in Table 7 of our experimental results, APS consistently improves document image tasks across a range of MLLMs with different resolutions. Additionally, results on TextVQA demonstrate that APS does not lead to performance degradation in tasks involving scene images.
The redundant regions in document images are often simple, such as plain backgrounds, distinguishing these regions is relatively easy. This raises the question of whether DTS’s effectiveness is due to this simplicity. Although this paper primarily explores VDU tasks, we are also interested in the generalizability of DTS. Therefore, we conducted some experiments on natural scene images. Specifically, we retrained a scene text image full-text recognition model using some open-source scene text data [10, 53, 50, 68], performing only the first pre-training phase. We then applied DTS to the trained model and tested it on the TextVQA dataset. Some visual results are shown in Fig. 5. It can be seen that DTS is effective in distinguishing nonessential regions even in the complex backgrounds of natural scene images.
5 Limitations
While DocKylin exhibits outstanding performance in visual document understanding (VDU) tasks, it currently lacks versatility for more general tasks, such as question answering and captioning in natural scenes. This limitation primarily stems from our current image encoder, which has not undergone extensive image-text pair pre-training and also possesses insufficient parameter capacity. Future work could explore scaling up the Swin Transformer and conducting pre-training similar to that of the current CLIP ViT. Alternatively, a dual-encoder approach [32, 70] could be employed, where a flexible, lightweight Swin Transformer is specifically used for extracting fine textual details within images, while a larger-capacity ViT model is utilized for capturing more general content. The combined image features from both encoders could then be integrated into the LLM.
6 Conclusions
@@
To address the issues of high input resolution and excessively long token sequences in current MLLMs when counters with VDU tasks, we propose the DocKylin model. It includes a parameter-free Adaptive Pixel Slimming (APS) preprocessing module that removes redundant regions from document images, effectively reducing their resolution. The model employs a flexible image encoding approach to adapt to the varying resolutions and aspect ratios of document image scenarios. Moreover, we propose a parameter-free token compression method, Dynamic Token Slimming (DTS), which clusters and selects essential tokens from a long visual sequence and employs aggregation to avoid potential information loss.
Experiments demonstrate that DocKylin achieves leading performance across multiple VDU benchmarks. APS and DTS effectively shorten the length of visual tokens, enhancing both computational efficiency and model performance. Additionally, we have validated the potential of applying these methods to other MLLMs or other scenarios beyond document images.
Appendix
Appendix A Benchmark Datasets Evaluation Metrics
@@
DocVQA [42], InfoVQA [41], ChartQA [40], WTQ [47], and VisualMRC [59] are all datasets designed for visual document image question answering. DocVQA encompasses a variety of document types, featuring a rich diversity of questions and data types. InfoVQA focuses on infographics that contain an abundance of visual elements. ChartQA and WTQ specifically target question answering for chart data and tabular data, respectively. VisualMRC, on the other hand, places a greater emphasis on developing natural language understanding and generation capabilities. DeepForm [57], FUNSD [19], SROIE [18], and POIE [25] are key information extraction datasets where the model is required to identify the value corresponding to a given key.
@@
Widely used text-centric benchmarks employ diverse metrics to better capture the distinct characteristics among various datasets. As depicted in Table 4 within the main text, both DocVQA [42] and InfoVQA [41] utilize ANLS [4] as their metric. The results reported are evaluated on their respective official websites. ChartQA [40] is evaluated by accuracy relaxed accuracy [43], while VisualMRC [59] relies on CIDEr [67]. DeepForm [57] is measured by F1 score, and WTQ [46] evaluated by accuracy.
Appendix B More Details and Results of Adaptive Pixel Slimming
@@
Appendix C More Details of DTS
Appendix D Declaration of Potential Harmful Content
While our model performs well on many tasks, we acknowledge that, like all MLLMs, due to the breadth of its training data and the diversity of its content, our model may occasionally generate inaccurate, misleading, or harmful content.
We hereby declare that, despite our efforts to mitigate this risk by selecting appropriate training data and adjusting model parameters, we cannot entirely eliminate the possibility of the MLLM producing harmful content. Therefore, when referring to this paper or using the model proposed herein, please be aware that the generated text may contain uncertainties and risks. We encourage readers to exercise caution and independently assess the accuracy and credibility of the content generated by the MLLM.

### v3

@@ Introduction
The pseudo-code for the specific implementation can be found in the supplementary materials.
@@
| FastV (K=2, R=50%) (Chen et al. 2024) | ✗ | 306 | 8.4 | 306 | 14.4 | 306 | 2.0 | |
| PruMerge (Shang et al. 2024) | ✓ | 123 | 5.3 | 123 | 13.8 | 121 | 0.4 | |
| PruMerge+ (Shang et al. 2024) | ✓ | 232 | 5.2 | 233 | 14.0 | 229 | 0.3 | |
@@
Experiments
Implementation Details
Our model underwent two training stages: pre-training and instruction tuning. All the training data are sourced from open-access datasets. Due to page limitations, specific datasets used are detailed in the supplementary materials. During the pre-training stage, we trained only the visual encoder and the MLP, keeping the LLM parameters fixed. In this phase, the model learns to perform full-text recognition on document images and convert charts and tables into markdown formats (Hu et al. 2024). The goal is to enable the visual encoder to comprehend document content and align the visual features with the language space through the MLP.
During the pre-training phase, the model was trained for 450,000 iterations with a batch size of 8. The learning rate decayed from 1e-4 to 1e-5 using cosine annealing, which took approximately 40 A800 GPU-days. In the instruction tuning phase, the model underwent 300,000 iterations with the same batch size and learning rate as in the pre-training phase. This phase further consumed approximately 24 A800 GPU-days. Our Adaptive Pixel Slimming method is employed in both phases, while the Dynamic Token Slimming (DTS) is used only during the instruction tuning phase. The maximum allowed number of pixels, MAX_SIZE, is set to 17281728. We initialize our visual encoder with Donut-Swin (0.07B) (Kim et al. 2022), pre-trained on the reading texts task, and use Qwen-7B-Chat (Bai et al. 2023a) as our LLM.
Comparisons with Prior Arts
We first compare our model with several existing general-purpose multimodal large language models (MLLMs). As these models sometimes tend to provide detailed responses, following (Zhang et al. 2024), we adopt accuracy as our metric: a response is considered correct if it contains the complete answer. The results, as shown in Table 1, demonstrate that DocKylin achieves a significant advantage over the competing models. This highlights that VDU tasks remain challenging for current general-purpose MLLMs. Note that descriptions of all these benchmarks are included in the supplementary materials.
We also compare our model with several existing text-centric MLLMs. The results, presented in Table 2, use the same metrics as those employed in the original benchmarks. Our method demonstrates advantages over these text-centric MLLMs, achieving leading performance across multiple benchmarks. Notably, our visual encoder is more lightweight than those used by the compared methods.
| APS | DTS | Token Len. | Train | Infer. | DocVQAval | ChartQA |
| ✗ | ✗ | 2770 | 1 | 1 | 63.1 | 55.8 |
| ✓ | ✗ | 2198 | 0.97 | 0.81 | 64.5 | 58.5 |
| ✓ | ✓ | 1250 | 0.86 | 0.61 | 65.1 | 59.0 |
@@
To further demonstrate the superiority of our method, we compared it with other parameter-free visual token compression methods, using LLaVA1.5 (Liu et al. 2023a) as our baseline model (without retraining). The results in Table 3 indicate that our APS and DTS achieve performance improvements with shorter sequence lengths compared to the baseline model, significantly outperforming existing SOTA compression methods. It is worth noting that these results are achieved without retraining. Typically, after retraining, models can better adapt to the new mode of compressed token sequences, leading to further performance improvements (Shang et al. 2024; Chen et al. 2024).
| Strategy | Scanned Document | Infographics | Chart | Table | Receipt | Ingredient List | ||||||
| Ratio | Acc. | Ratio | Acc. | Ratio | Acc. | Ratio | Acc. | Ratio | Acc. | Ratio | Acc. | |
| None | - | 55.2 | - | 26.7 | - | 73.5 | - | 34.8 | - | 44.0 | - | 25.9 |
| APS-mask | 36.6 | 54.7 | 16.7 | 26.4 | 31.5 | 74.8 | 4.0 | 34.6 | 63.9 | 44.0 | 0.6 | 26.1 |
| DTS-mask | 48.2 | 54.9 | 35.2 | 26.3 | 42.4 | 73.8 | 38.1 | 31.5 | 44.1 | 43.2 | 34.0 | 25.2 |
| DTS-R-mask | 51.8 | 0.2 | 64.8 | 14.6 | 57.6 | 4.4 | 61.9 | 0.6 | 55.9 | 0.1 | 66.0 | 0.5 |
Ablation Studies
@@
We validate the effectiveness of APS and DTS on DocKylin and present the results in Table 4. Note that, settings without DTS in Table 4 refer to retraining a new model from scratch without employing DTS in the instruction tuning phase, rather than simply removing DTS from the already trained DocKylin model. The results demonstrate that both APS and DTS effectively reduce token sequence length and enhance computational efficiency, while also improving performance. This aligns with findings from previous studies (Liu et al. 2024b), which suggest that excessively long sequences due to high resolution may degrade performance. This degradation likely occurs because the visual sequences contain too many redundant tokens, placing greater demands on the model’s ability to handle long contexts and making the learning more challenging. Furthermore, the APS and DTS methods can also be used during the training phase to accelerate training. Specifically, applying the DTS method during the instruction tuning phase can reduce training time by approximately 15%.
| Methods | Resolution | DocVQAval | InfoVQAval | SROIE | FUNSD |
| LLaVA1.5 (Liu et al. 2023a) | 336336 | 8.5 | 14.7 | 1.7 | 0.2 |
| LLaVA1.5 (Liu et al. 2023a) + APS | 10.7 (+27.4%) | 14.7 (+0%) | 3.7 (+118%) | 0.92 (+360%) | |
| Qwen-VL (Bai et al. 2023b) | 448448 | 48.1 | 23.9 | 34.5 | 20.6 |
| Qwen-VL (Bai et al. 2023b) + APS | 51.2 (+6.4%) | 24.7 (+4.1%) | 40.0 (+15.9%) | 24.3 (+17.9%) | |
| Monkey (Li et al. 2023b) | 896896 | 50.1 | 25.8 | 41.9 | 24.1 |
| Monkey (Li et al. 2023b) + APS | 56.3 (+12.4%) | 27.5 (+6.6%) | 47.0 (+12.2%) | 27.3 (+13.3%) |
We conduct experiments to further explore: 1) How accurately do APS and DTS identify redundant regions? 2) Do APS and DTS demonstrate consistent performance across different data types? Specifically, we apply APS and DTS to identify redundant regions in different document images, and then mask these regions instead of removing them (as shown in Fig. 5). The masked images are used as inputs to an existing MLLM to assess any performance degradation. This approach allows us to isolate the accuracy of APS and DTS from the performance improvements brought by redundancy removal.
Table 5 shows that APS effectively reduces resolution across all image types with minimal performance impact. However, the reduction effects on the table and ingredient list are less pronounced, likely due to the lines spanning the entire image, making APS difficult to identify redundant rows or columns. DTS also effectively reduces resolution across all image types and typically achieves greater resolution reduction than APS, as it can identify local areas beyond entire rows and columns. While DTS may cause noticeable performance drops in the table datasets, our visual analysis suggests this is due to new patterns introduced by masking table lines, not the masking of critical text content. Practically, such potential performance degradation from DTS can be largely mitigated: 1) DTS retains information by similarity-weighted aggregation approach rather than completely discarding redundant tokens; and 2) during image encoding, information from masked areas is integrated into neighboring tokens.
We also conduct additional experiments where we reversed the DTS-mask results by masking the non-redundant regions identified by DTS, denoted as DTS-R-mask. The results of the DTS-R-mask indicate a significant performance drop, further demonstrating the accuracy of DTS in identifying redundant regions.
@@
Since APS is a parameter-free method that requires no additional training or modification of model architecture, it can be easily applied to other MLLMs to enhance their performance on high-resolution document images. To further demonstrate its effectiveness, we apply APS to various existing MLLM models (employing it before the image processing module of these models). As shown in Table 6, APS consistently improves document image tasks across a range of MLLMs with different resolutions.
Conclusions
@@
To address the issues of high input resolution and excessively long token sequences in current MLLMs when counters with VDU tasks, we propose the DocKylin model. It includes an Adaptive Pixel Slimming (APS) preprocessing module that removes redundant regions from document images, effectively reducing their resolution. The model employs a lightweight and flexible image encoding approach to adapt to the varying resolutions and aspect ratios of document image scenarios. Moreover, we propose a token compression module, Dynamic Token Slimming (DTS), which clusters and selects essential tokens from a long visual sequence and employs aggregation to avoid potential information loss. Experiments demonstrate that DocKylin achieves leading performance across multiple VDU benchmarks. Extensive experiments demonstrate that APS and DTS effectively reduce the length of visual tokens, enhancing both computational efficiency and model performance. The parameter-free nature of our APS and DTS modules not only enhances efficiency but also facilitates easy integration into existing multimodal large language models. This feature makes our approach particularly valuable for improving document understanding capabilities in a wide range of applications.
Appendix A Appendix
Training Data
| Stage | Datasets |
| Pre-training | CASIA-HWDB (Liu et al. 2011), CCpdf (Turski et al. 2023), ChartQA (Masry et al. 2022), DeepForm (Svetlichnaya 2020), DocVQA (Mathew, Karatzas, and Jawahar 2021), DVQA (Kafle et al. 2018), FigureQA (Kahou et al. 2017), InfoVQA (Mathew et al. 2022), KLC (Stanisławek et al. 2021), OCRVQA (Mishra et al. 2019), PlotQA (Methani et al. 2020), PubTabNet (Zhong, ShafieiBavani, and Jimeno Yepes 2020), RVL-CDIP (Harley, Ufkes, and Derpanis 2015), SCUT-HCCDoc (Zhang, Liang, and Jin 2020), SynthDog (Kim et al. 2022), TAT-QA (Zhu et al. 2021), VisualMRC (Tanaka, Nishida, and Yoshida 2021), WildReceipt (Sun et al. 2021) |
| Instruction tuning | ChartQA (Masry et al. 2022), DeepForm (Svetlichnaya 2020), DocILE (Šimsa et al. 2023), DocVQA (Mathew, Karatzas, and Jawahar 2021), DVQA (Kafle et al. 2018), InfoVQA (Mathew et al. 2022), InsuranceQA (Studio 2022), KLC (Stanisławek et al. 2021), OCR-TROSD(Tianchi 2020), TAT-QA (Zhu et al. 2021), VisualMRC (Tanaka, Nishida, and Yoshida 2021), WildReceipt (Sun et al. 2021), WTQ (Pasupat and Liang 2015a) |
The training data for the two stages are summarized in Table 7.
| Methods | Resolution | TextVQA |
| LLaVA1.5 (Liu et al. 2023a) | 336336 | 38.7 |
| LLaVA1.5 (Liu et al. 2023a) + APS | 40.2 (+4.1%) | |
| Qwen-VL (Bai et al. 2023b) | 448448 | 62.1 |
| Qwen-VL (Bai et al. 2023b) + APS | 62.4 (+0.4%) | |
| Monkey (Li et al. 2023b) | 896896 | 64.3 |
| Monkey (Li et al. 2023b) + APS | 65.2 (+1.4%) |
Evaluation Metrics for Each Benchmark
@@ Appendix A Appendix
DocVQA (Mathew, Karatzas, and Jawahar 2021), InfoVQA (Mathew et al. 2022), ChartQA (Masry et al. 2022), WTQ (Pasupat and Liang 2015b), and VisualMRC (Tanaka, Nishida, and Yoshida 2021) are all datasets designed for visual document image question answering. DocVQA encompasses a variety of document types, featuring a rich diversity of questions and data types. InfoVQA focuses on infographics that contain an abundance of visual elements. ChartQA and WTQ specifically target question answering for chart data and tabular data, respectively. VisualMRC, on the other hand, places a greater emphasis on developing natural language understanding and generation capabilities. DeepForm (Svetlichnaya 2020), FUNSD (Jaume, Ekenel, and Thiran 2019), SROIE (Huang et al. 2019), and POIE (Kuang et al. 2023) are key information extraction datasets where the model is required to identify the value corresponding to a given key.
@@
Widely used text-centric benchmarks employ diverse metrics to better capture the distinct characteristics among various datasets. As depicted in Table 2 within the main text, both DocVQA (Mathew, Karatzas, and Jawahar 2021) and InfoVQA (Mathew et al. 2022) utilize ANLS (Biten et al. 2019) as their metric. The results reported are evaluated on their respective official websites. ChartQA (Masry et al. 2022) is evaluated by accuracy relaxed accuracy (Methani et al. 2020), while VisualMRC (Tanaka, Nishida, and Yoshida 2021) relies on CIDEr (Vedantam, Lawrence Zitnick, and Parikh 2015). DeepForm (Svetlichnaya 2020) is measured by F1 score, and WTQ (Pasupat and Liang 2015a) is evaluated by accuracy.
Potential Applications for Scene Text Images
APS is primarily designed for document images, where redundant regions are commonly distinctly defined. In contrast, elements in natural scene images tend to be more complex, making it challenging to identify extensive redundant regions. Consequently, APS generally has a limited impact on natural scene tasks. To validate this, we applied APS to the TextVQA dataset. As shown in Table 8, APS does not introduce any negative impact. Some visualization results of APS in this scenario can be found in the last row of Fig. 7.
The redundant regions in document images are often simple, such as plain backgrounds, distinguishing these regions is relatively easy. This raises the question of whether DTS’s effectiveness is due to this simplicity. Although this paper primarily explores VDU tasks, we are also interested in the generalizability of DTS. Therefore, we conducted some experiments on natural scene images. Specifically, we retrained a scene text image full-text recognition model using some open-source scene text data (Ch’ng and Chan 2017; Singh et al. 2019; Shi et al. 2017; Veit et al. 2016), performing only the first pre-training phase. We then applied DTS to the trained model and tested it on the TextVQA dataset. Some visual results are shown in Fig. 6. It can be seen that DTS is effective in distinguishing nonessential regions even in the complex backgrounds of natural scene images.
More Details and Results of Adaptive Pixel Slimming
@@
More Details and Results of Dynamic Token Slimming
The pseudocode for the Similarity Weighted Aggregation method in DTS is shown in Algorithm 3.
We present the visualization results of APS and DTS on different data types in Figure 1. From these visualizations, we can draw the following conclusions: 1) APS effectively removes redundant regions when document backgrounds are simple, but struggles in scenarios involving graphics, document borders, table lines, or photographs. 2) In contrast, DTS is more robust and can handle these challenging cases that APS cannot. 3) While DTS is not perfect and may mistakenly classify some text areas as redundant, the impact of this issue in practical applications is minimal because: 1) As detailed in the methodology section, DTS does not completely discard redundant tokens but retains as much information as possible through similarity-weighted aggregation; 2) Although DTS might mask useful areas at the image level, during the image encoding process, information from these regions is integrated into neighboring tokens.